OpenStudy (anonymous):

In the formula for the volume of a cylinder, how is the area of the base determined?

OpenStudy (juicstice):

by finding the area of the circle

OpenStudy (anonymous):

my choices are \[A.\pi \] \[B.r^{2}\] \[C.\pi r^{2}\] \[D.2\pi \] @Juicstice

OpenStudy (juicstice):

\[\Pi \times r ^{2}\]

OpenStudy (juicstice):

that's how you find the area of a circle

OpenStudy (anonymous):

thank you